# Spectroscopy and Coherence of an Excited-State Transition in Tm3+: YAlO3 at Telecommunication Wavelength

### v1

@@ Abstract
We characterize the spectroscopic and coherence properties of the 1451.37 nm excited-state zero-phonon line (ZPL) between the 3F4 and the 3H4 manifolds of a thulium-doped yttrium aluminum perovskite (Tm:YAlO3) crystal at temperatures around 1.5 K. We measure the absorption spectrum between the3H6 - 3F4 and 3F4 - 3H4 manifolds, the inhomogeneous broadening of the 3F4 - 3H4 (excited-state) ZPL, and the lifetimes of the higher-lying and lower-lying excited states. We also investigate spectral hole-burning spectra with varying magnetic fields, which provides insight into hyperfine and superhyperfine interactions. Using again spectral holes but also optical free induction decays (FIDs), we assess optical coherence times, finding a maximum of 1.1 $\mu s$ when a 2T magnetic field is applied. Our results—the first to demonstrate coherence of an excited-state transition in a rare-earth crystal—suggest the possibility of exploiting such transitions for quantum technology.
@@ Acknowledgements
The authors thank Charles Thiel, Rufus Cone and their team for early input to these studies, and Javier Carrasco Ávila, Archi Gupta and Deeksha Gupta for discussions. We acknowledge funding through the Netherlands Organization for Scientific Research, the European Union’s Horizon 2020 Research and Innovation Program under Grant Agreement No. 820445 and Project Name Quantum Internet Alliance, and the Swiss State Secretariat for Education, Research and Innovation (SERI) under Contract Number UeM019-3.

### v2

@@ Abstract
We characterize the spectroscopic and coherence properties of the 1451.37 nm excited-state zero-phonon line (ZPL) between the 3F4 and the 3H4 manifolds of a thulium-doped yttrium aluminum perovskite (Tm3+:YAlO3) crystal at temperatures around 1.5 K. We measure the absorption spectrum between the3H6 - 3F4 and 3F4 - 3H4 manifolds, the inhomogeneous broadening of the 3F4 - 3H4 (excited-state) ZPL, and the lifetimes of the higher-lying and lower-lying excited states. We also investigate spectral hole-burning spectra with varying magnetic fields, which provides insight into hyperfine and superhyperfine interactions. Using again spectral holes but also optical free induction decays (FIDs), we assess optical coherence times, finding a maximum of 1.1 $\mu s$ when a 2T magnetic field is applied. Our results—the first to demonstrate coherence of an excited-state transition in a rare-earth crystal—suggest the possibility of exploiting such transitions for quantum technology.
@@ Acknowledgements
The authors thank Javier Carrasco Ávila, Archi Gupta and Deeksha Gupta for discussions. WT acknowledges funding through the Netherlands Organization for Scientific Research, the European Union’s Horizon 2020 Research and Innovation Program under Grant Agreement No. 820445 and Project Name Quantum Internet Alliance, and the Swiss State Secretariat for Education, Research and Innovation (SERI) under Contract Number UeM019-3. Work at MSU was supported in part by the Air Force Research Laboratory under award number FA8750-23-2-0500.
Data availability: The data that support the findings of this article are openly available [li_2025_17553877].